1)Doctors order is Demerol 5mg/kg PO q4h pm pain. The patient weighs 200lbs. Demerol comes in 25mg/ml. How many ml would you give?

2)Prednisone 10 mg/kg/day IV every 12 hrs is ordered.On hand you have Prednisone 50mg/ml. The child weighs 125kg. a)how many ml per dose will you give? b)how many mg per dose will the patient receive? c) how many mg per day will the patient receive?

3)Augmentin 10 mg/kg PO every 12 hours is ordered. When Augmentin is reconstituted it yields 125mg/ml. The patient weighs 200ibs. How many ml will you give?

4)You must administer a Heparin bolus prior to starting a gtt. You have Heparin vials marked 5000 units/ml. You must administer 1000 units of Heparin IV, how many ml do you administer?
Answer #1

1.)

Step 1: Convert pounds into kilogram

The patient weighs 200lbs.

200 x 0.453 = 90.7kg.

Step 2: Calculate the desired dose

The physician order is 5mg/kg.

So for 90.7 kg = 90.7 x 5 = 453.5 mg.

Step 3: Calculate the actual dose

Actual Dose = Desired Dose / Dose on Hand x Quantity

Desired Dose = 453.5mg

Dose on Hand = 25mg/mL

Quantity = 1mL

Actual Dose = 453.5 / 25 x 1

= 18.14mL

Demerol 18.1mL to be given.

2.)

Step 1: calculate the ordered dose

Given order 10mg/kg/day.

the patient weighs 125kg.

Therefore the desired dose is 125 x 10 = 1250mg/ day.

Step 2: Dose on hand

Prednisolone 50mg/mL is available.

Step 3:

a)how many ml per dose will you give?

Total dose for a day = 1250 mg.

Dose ordered every 12 hours.

So single dose = 1250/2 =625mg.

Total mL per dose = Desired Dose / Dose on Hand x Quantity

= 625 / 50 x 1

= 12.5mL

For per dose 12.5mL to be given.

b)how many mg per dose will the patient receive?

Total mg per day = 1250mg / day for every 12hours.

Total mg per dose = 1250/2 = 625mg.

c) how many mg per day will the patient receive?

Total mg per day = 1250mg.

3.)

Step 1: Convert weight to kg.

Patient weighs 200lbs.

200 x 0.453 = 90.7 kg.

Step 2: Calculate the ordered dose.

The ordered dose is 10mg/kg.

The patient weighs 90.7 kg  

90.7 x 10 = 907mg.

Step 3: Calculate the actual dose:

Actual dose = Desired dose / Dose on hand x quantity

= 907 / 125 x 1

= 7.25mL

So, Augmentin 7.25mL to be administered.

4.) Calculate Heparin dosage

Desired dose = 1000 units

Dose on Hand = 5000units.

Quantity = 1mL

Dose = Desired dose / Dose on hand x Q

= 1000 / 5000 x 1

= 0.2mL

Hence 0.2 mL of Heparin bolus should be administered.